OpenStudy (anonymous):

Gallium has a weighted average atomic mass of 69.7 and two naturally occurring isotopes: gallium-69 and gallium-71. Gallium-69 has a relative abundance of 60.1% and a mass of 68.9. The relative abundance of gallium-71 is 39.9%. Based on this information, what is the mass of gallium-71? 71.1 g/mol 71.1 g/mol 70.9 g/mol 70.9 g/mol

OpenStudy (somy):

what have u done so far?

OpenStudy (anonymous):

i was hoping you could help me start

OpenStudy (anonymous):

This is a pretest so idk too much of this stuff

OpenStudy (somy):

sorry its lagging so bad

OpenStudy (anonymous):

No worries

OpenStudy (somy):

okay Gallium has a weighted average atomic mass of 69.7 isotopes: Gallium-69 has a relative abundance of 60.1% and a mass of 68.9. The relative abundance of gallium-71 is 39.9%. Based on this information, what is the mass of gallium-71? so the formula is like this \[\frac{ \sf (PAOI_1 \times mass~of~isotope_1) \times (PAOI_2 \times mass~of~isotope_2) }{ 100}\sf =average~ atomic~ mass\] PAOI = percentage abundance of isotope

OpenStudy (somy):

you have isotope 2 isotopes so i made the formula according to it so u have isotope number 1(Gallium-69) and isotope number 2 (gallium-71) isotope number 1 has percentage abundance and mass isotope number 2 has percentage abundance given but MASS is missing so is unknown X and you are given average atomic mass which is 69.7

OpenStudy (somy):

use the formula and find unknown X mass of gallium-71

OpenStudy (somy):

oh wait sorry my mistake there!

OpenStudy (somy):

\(\sf \frac{(PAOI_1 \times mass~of~isotope_1) + (PAOI_2 \times mass~of~isotope_2)}{100}= average~ atomic~ mass\)

OpenStudy (somy):

corrected

OpenStudy (anonymous):

its 70.9

OpenStudy (anonymous):

right?

OpenStudy (somy):

sec let me calculate lol

OpenStudy (anonymous):

And when you're done can yu help me with one more please?

OpenStudy (somy):

yep u r correct :3

OpenStudy (somy):

yep sure as long as i can i will help

OpenStudy (anonymous):

The elements located in the first column, or group, in the periodic table all share a similar electron configuration. If the atom is neutral, how many electrons are found in the outermost shell of each of these elements?

OpenStudy (anonymous):

1 2 3 4

OpenStudy (somy):

neutral means 0 charge what do u think?

OpenStudy (anonymous):

is it 1?

OpenStudy (somy):

how many electrons do group 1 elements have in their outermost shell? that is when the atom is in neutral state when atom loses or gains electrons -in other words its electron number on outermost shell increases or decreases from its original number of electrons - is when the atom becomes charged

OpenStudy (somy):

so yeah bingo!

OpenStudy (anonymous):

You're the best, thanks ! (:

OpenStudy (somy):

no problem :))